Problem 1. The following probability table is a breakdown of managers in four sectors of eco omy: Gender Banking Insurance Real Estate Manufacturing Male12 Feal.08 .14 ,13 .21 .06 ,17 .09 a. What is the probability that the selected manager is female? b. What is the probability that the selected manager is female or in insurance business? c. What is the probability that the selected manager is neither female nor in insurance business? d. Suppose the selected manager is female. What is the probability that she is real estate business? e. Is the event being female independent of the event of begin in insurance business? explai

Please include the process of the answer.
Answer #1

(a) P(Female) = 0.08 + 0.06 + 0.17 + 0.09 = 0.40

(b) P(Female or Insurance) = P(Female) + P(Insurance) - P(Female and Insurance) = 0.40 + (0.14 + 0.06) - 0.06 = 0.54

(c) P(Neither Female nor Insurance) = 1 - P(Female or Insurance) = 1 - 0.54 = 0.46

(d) P(Real Estate | Female) = P(Real Estate and Female) / P(Female) = 0.17/0.40 = 0.425

(e) P(Female | Insurance) = P(Female and Insurance) / P(Female) = 0.06/0.40 = 0.15

P(Female) = 0.40

Since 0.15 ≠ 0.40. the events Female and Insurance are not independent.